OpenStudy (sgayton27):

Find the domain. PLEASE HELP! g(x)=sqrt (6+x)/7-x

OpenStudy (sgayton27):

\[g(x)=\frac{ \sqrt{6+x} }{ 7-x }\]

OpenStudy (anonymous):

\[x \neq7 , x \ge 6\]